The joint age of uncle Tom and his brother is 60 years. In 6 years time uncle Tom will be twice as old as his brother was 6 years ago.How old is Uncle Tom

Answer by ankor@dixie-net.com(22740)

t = Tom's age
b = brother's age
Write an equation for each statement
:
The joint age of uncle Tom and his brother is 60 years.
t + b = 60
b = (60-t), use this form for substitution
In 6 years time uncle Tom will be twice as old as his brother was 6 years ago.
t + 6 = 2(b - 6)
t + 6 = 2b - 12
t = 2b - 12 - 6
t = 2b - 18
How old is Uncle Tom
In the above equation, replace b with (60-b)
t = 2(60-b) - 18
t = 120 - 2b - 18
t + 2b = 120 - 18
3t = 102
t = 102/3
t = 34 yrs is Tom's age